# {dedication}

### v1

@@ 4 Benchmark Suite Design
Abbv
GNN Model
Application Domain
Graph Input Type
PSAGE
PinSAGE
Recommendation
Heterogeneous Graph
STGCN
Spatio Temporal GCN
Traffic Forecasting
Dynamic Graph
DGCN
Deep GCN
Molecular Property Prediction
Homogeneous Graph
GW
GraphWriter
Text Generation
Heterogeneous Graph
KGNN
k Graph Neural Networks
Protein Classification
Homogeneous Graph
ARGA
Adverserially Regularized Graph Autoencoder
Node Clustering
Homogeneous Graph
TLSTM
Tree Long Short-Term Memory Networks
Sentiment Classification
Homogeneous Graph
@@
Abbv
Datasets
# Node # Edge
PSAGE
Nowplaying (NWP) [169]
22.9M 1.9M
Movielens (MVL) [64]
1.9M 9.7K
STGCN
LA [168]
207 325
PEMS_Bay (PEMS) [168]
1722 2694
DGCN
MOLHIV [72]
1.04M 1.1M
MOLTOX [72]
145K 151K
GW
AGENDA  [87]
885K 2.57M
KGNN
Proteins (PROT) [80]
43K 162K
ARGA
Cora [163]
2K 10.5K
CiteSeer (CSEER) [163]
3.3K 9.2K
PubMed (CSEER) [163]
19.7K 88.6K
TLSTM
Stanford Sentiment Treebank (SNTM) [131]
318K 310K
@@
k-GNNs: Most GNN models are one-dimensional in nature and cannot effectively capture any higher-order information, such as the properties of subgraphs, within the graph. As a result, they fail the graph isomorphism test proposed by Weisfeiler and Lehman [151] (WL algorithm). The WL algorithm is a test used to determine the expressiveness power of a GNN by testing if an algorithm is able to distinguish whether two graphs are isomorphic or not. Two graphs are said to be isomorphic if they have the same number of vertices, edges, and connectivity. Therefore, researchers have developed higher-dimensional hierarchical GNNs, called k-GNNs (where the k stands for the dimension), which can capture properties of subgraphs [104]. This enables GNNs to perform close to the k-WL graph isomorphism test [104]. We include two variants of k-GNNs, (KGNNL and KGNNH to denote a lower and higher dimensional version of k-GNN, respectively) and use them to perform classification of protein molecules. The primary reason we include this workload in our suite is to study how application characteristics and behavior change as we move towards higher-dimension GNNs [118].
@@
Tree-LSTM: Sentiment classification is an important task in the Natural Language Processing (NLP) domain. Tree Long Short-Term Memory Networks (Tree-LSTMs) [138] are one group of models developed for this task. In contrast to the linear model used in an LSTM, Tree-LSTMs use a tree-structured network topology and can outperform linear LSTMs in the sentiment classification task [138]. The Tree-LSTM method implemented in DGL uses the idea of batching. The basic idea of batching is to collect smaller graphs that are part of the dataset and convert them into a batched larger graph. We include the Tree-LSTM model in GNNMark to study how batching multiple small graphs to a larger graph impacts the behavior of an application.
@@ 7 GNNMark Summary
In this thesis proposal , we present GNNMark, a diverse benchmark suite of GNN workloads designed for the characterization of GPU performance. To the best of our knowledge, we are the first to propose a GNN training focused benchmark suite for the architecture community. We use GNNMark to perform a detailed characterization of GNNs to understand the architectural implications of training on GPU systems. Our work provides novel insights that show the major architectural bottlenecks in GNN training and suggests how they can be potentially addressed.
@@
Chapter 4 Acceleration of GNNs
Graph Neural Networks (GNNs) are characterized by their computational structure, which involves sparse matrix operations in the combination phase and dense operations in the aggregation phase. This section focuses on the acceleration of the combination phase through the development of an accelerator specifically designed for Sparse General Matrix-Matrix Multiplication (SpGEMM). Subsequently, we introduce an accelerator designed to efficiently manage both the sparse (combination) and dense (aggregation) computational phases of GNN workloads, thereby establishing a comprehensive and versatile GNN accelerator.
@@
As part of this thesis proposal , we explore a novel distributed memory SpGEMM implementation. We specifically target this work for a custom accelerator. Our approach improves performance by mapping the computation of row-wise products, hashtables, and on-chip accumulation to the accelerator’s scratchpads. We provide a new set of performance metrics for this class of workload and demonstrate their utility using a suite of micro-benchmarks run using synthetic, as well as real-world, datasets. We also introduce a new Memory-aware Aligned Parallel Compressed Sparse Row matrix storage format called MAP-CSR to further accelerate local memory accesses. Running on a custom graph-based accelerator, we are able to achieve consistent speedup over MKL and provide insights on the scalability of our implementation.
@@ 1 Motivation for SpGEMM kernel acceleration
    We characterize the challenges faced while developing efficient SpGEMM kernels. To this end, we perform an analysis of various sparse matrix multiplication methods and evaluate the advantages and limitations of each.
@@ 2 Background on SpGEMM
Variations in the implementations of these two phases give rise to various SpGEMM algorithms. There are four methods to compute the first multiplication phase, as shown in Figure 3. Each method exhibits different memory access patterns and provides varying degrees of parallelism.
@@ 3 MAP - CSR Storage Format
The Conventional CSR format allows only sequential write operations, which poses a considerable challenge to implement SpGEMM kernels that scale on multi-node systems. It also introduces many synchronization operations, leading to performance degradation. We introduce a novel matrix storage format called MAP-CSR, that is designed to scale well on large-scale distributed systems.
@@ 4 SMASH Kernel
Our novel implementation of the SpGEMM kernel is based on a row-wise product method called SMASH. Our method exploits high data reuse behavior [120] and minimizes the number of input matrix reads, while still maintaining low on-chip memory usage. SMASH incorporates on-chip memory to store intermediate results and leverages the atomic instructions to accumulate these partial products.
In this thesis proposal , we present a set of successive improvements, resulting in three versions of SMASH. In each version we identify the remaining bottlenecks, and then optimize our algorithm to mitigate them in the next version. Each SMASH implementation targets a specific performance bottleneck on the custom accelerator architecture.
The following subsections describe our implementation of SMASH, along with three different optimizations. Similar to Gustafson’s two-phase matrix multiplication algorithm [62], our SMASH implementation is characterized by two phases:
@@ 1 Memory computation phase
Analogous to Gustafson’s two-phase matrix multiplication approach [62], the first phase of SMASH determines the memory required for the output matrix $C$, as well as the on-chip memory requirements of the intermediate products. For evaluation purposes, we do not include the time consumed in the memory computation phase while considering speedup over other architectures. Evaluation methodologies are further discussed in Section 5. This phase can be further decomposed into three tasks.
@@
    Compute Memory Requirements - After reading the input matrix arrays in CSR format, we compute the required size of memory required to store the output matrix by counting the total FMA operations per row. We compute the maximum number of non-zeros for every row of output matrix $C$ using Gustafson’s two-step algorithm [62]. In this thesis proposal , we refer to this term as Floating-Point Multiply-Add (FMA). The computation of FMA s per row has a computational complexity of $\mathcal{O}(n)$, where $n$ is the size of the input matrix.
@@ 2 Product Computation Phase
1 // READ PHASE while Till you reach end of window do
2       // Atomically distribute work to each thread $token\leftarrow$ Each thread will receive one unique token if $token\_id\ \%\ 2=0$ then
@@ 5             $row\_begin\leftarrow A\_col\_ptr\_copy\_2[\frac{token\_id}{2}]$
7      $row\_end\leftarrow A\_col\_ptr[\frac{token\_id}{2}+1]$ for $i\leftarrow$ Iterate from $row\_begin$ to $row\_end$ do
@@
9                   $col\_begin\leftarrow B\_row\_ptr[\frac{token\_id}{2}]$ $col\_end\leftarrow B\_row\_ptr[\frac{token\_id}{2}+1]$ if $token\_id\ \%\ 2=0$ then
@@
2       // Multiply element from $mat\_A$ with that from $mat\_B$ and store its tag and value $tag\leftarrow$ $X$ coordinate from $mat\_A$ element and $Y$ coordinate from $mat\_B$ element // Hash the Tag $tag\leftarrow tag\ \%\ prime\_modulo$ if $SPAD\_tag[tag]=EMPTY$ then
3             $SPAD\_tag[tag]\leftarrow tag$ // Store Tag on scratchpad $SPAD\_val[tag]\leftarrow value$ // Store Value on scratchpad
@@ 5 Evaluation
We evaluate the performance of our SMASH SpGEMM kernel implementation on synthetic, as well as real-world, datasets. For synthetic datasets, we chose the RMAT, as it produces datasets possessing a non-zero power-law distribution [29], making it harder to find patterns in the non-zero values.
@@ 6 SMASH Summary
Chapter 5 Summary and Plans
The primary objective of this thesis is to provide a comprehensive understanding of the emerging domain of graph computing. It focuses on identifying and addressing the challenges in this field, with an emphasis on developing strategies for accelerating graph computing processes. The thesis aims to contribute significantly to the knowledge base in this area, offering innovative solutions and methodologies.
1 Work Completed to Date
GNNMark Benchmark Suite: The thesis introduces GNNMark, a diverse benchmark suite tailored for evaluating GPU performance in Graph Neural Network (GNN) workloads. This suite has facilitated a deeper understanding of inhereent characteristics specific to GNN training. Key findings from this work include:
@@ 1 Work Completed to Date
    Identification of major architectural bottlenecks in GNN training.
@@
    Proposals for potential mitigation strategies.

### v2

@@ 4 Benchmark Suite Design
Abbv GNN Model Application Domain Graph Input Type
PSAGE PinSAGE Recommendation Heterogeneous Graph
STGCN Spatio Temporal GCN Traffic Forecasting Dynamic Graph
DGCN Deep GCN Molecular Property Prediction Homogeneous Graph
GW GraphWriter Text Generation Heterogeneous Graph
KGNN k Graph Neural Networks Protein Classification Homogeneous Graph
ARGA Adverserially Regularized Graph Autoencoder Node Clustering Homogeneous Graph
TLSTM Tree Long Short-Term Memory Networks Sentiment Classification Homogeneous Graph
@@
Abbv Datasets # Node # Edge
PSAGE Nowplaying (NWP) [210] 22.9M 1.9M
Movielens (MVL) [75] 1.9M 9.7K
STGCN LA [209] 207 325
PEMS_Bay (PEMS) [209] 1722 2694
DGCN MOLHIV [83] 1.04M 1.1M
MOLTOX [83] 145K 151K
GW AGENDA  [105] 885K 2.57M
KGNN Proteins (PROT) [96] 43K 162K
ARGA Cora [203] 2K 10.5K
CiteSeer (CSEER) [203] 3.3K 9.2K
PubMed (CSEER) [203] 19.7K 88.6K
TLSTM Stanford Sentiment Treebank (SNTM) [165] 318K 310K
@@
k-GNNs: Most GNN models are one-dimensional in nature and cannot effectively capture any higher-order information, such as the properties of subgraphs, within the graph. As a result, they fail the graph isomorphism test proposed by Weisfeiler and Lehman [190] (WL algorithm). The WL algorithm is a test used to determine the expressiveness power of a GNN by testing if an algorithm is able to distinguish whether two graphs are isomorphic or not. Two graphs are said to be isomorphic if they have the same number of vertices, edges, and connectivity. Therefore, researchers have developed higher-dimensional hierarchical GNNs, called k-GNNs (where the k stands for the dimension), which can capture properties of subgraphs [129]. This enables GNNs to perform close to the k-WL graph isomorphism test [129]. We include two variants of k-GNNs, (KGNNL and KGNNH to denote a lower and higher dimensional version of k-GNN, respectively) and use them to perform classification of protein molecules. The primary reason we include this workload in our suite is to study how application characteristics and behavior change as we move towards higher-dimension GNNs.
@@
Tree-LSTM: Sentiment classification is an important task in the Natural Language Processing (NLP) domain. Tree Long Short-Term Memory Networks (Tree-LSTMs) [174] are one group of models developed for this task. In contrast to the linear model used in an LSTM, Tree-LSTMs use a tree-structured network topology and can outperform linear LSTMs in the sentiment classification task [174, 87]. The Tree-LSTM method implemented in DGL uses the idea of batching. The basic idea of batching is to collect smaller graphs that are part of the dataset and convert them into a batched larger graph. We include the Tree-LSTM model in GNNMark to study how batching multiple small graphs to a larger graph impacts the behavior of an application.
@@ 7 GNNMark Summary
In this dissertation , we present GNNMark, a diverse benchmark suite of GNN workloads designed for the characterization of GPU performance. To the best of our knowledge, we are the first to propose a GNN training focused benchmark suite for the architecture community. We use GNNMark to perform a detailed characterization of GNNs to understand the architectural implications of training on GPU systems. Our work provides novel insights that show the major architectural bottlenecks in GNN training and suggests how they can be potentially addressed.
@@
Chapter 4 Algorithmic Strategies for GNN Acceleration
Graph Neural Networks (GNNs) are characterized by their computational structure, which involves dense matrix operations in the combination phase and sparse matrix operations in the aggregation phase. This chapter focuses on the acceleration of the combination phase through the development of an accelerator specifically designed for Sparse General Matrix-Matrix Multiplication (SpGEMM). Subsequently, in the following chapter, we introduce an accelerator designed to efficiently manage both the sparse (combination) and dense (aggregation) computational phases of GNN workloads, thereby establishing a comprehensive and versatile GNN accelerator.
@@
As part of this dissertation , we explore a novel distributed memory SpGEMM implementation. We specifically target this work for a custom accelerator. Our approach improves performance by mapping the computation of row-wise products, hashtables, and on-chip accumulation to the accelerator’s scratchpads. We provide a new set of performance metrics for this class of workload and demonstrate their utility using a suite of micro-benchmarks run using synthetic, as well as real-world, datasets. We also introduce a new Memory-aware Aligned Parallel Compressed Sparse Row matrix storage format called MAP-CSR to further accelerate local memory accesses. Running on a custom graph-based accelerator, we are able to achieve consistent speedup over MKL and provide insights on the scalability of our implementation.
@@ 1 Motivation for SpGEMM kernel acceleration
    We characterize the challenges faced while developing efficient SpGEMM kernels. To this end, we perform an analysis of various sparse matrix multiplication methods and evaluate the advantages and limitations of each..
@@ 2 Background on SpGEMM
Variations in the implementations of these two phases give rise to various SpGEMM algorithms. There are four methods to compute the first multiplication phase, as shown in Figure 3 and Figure 2. Each method exhibits different memory access patterns and provides varying degrees of parallelism.
@@ 3 MAP - CSR Storage Format
The Conventional CSR format allows only sequential write operations, which poses a considerable challenge to implement SpGEMM kernels that scale on multi-node systems. It also introduces many synchronization operations, leading to performance degradation. We introduce a novel matrix storage format called MAP-CSR [156], that is designed to scale well on large-scale distributed systems.
@@ 4 SMASH Kernel
Our novel implementation of the SpGEMM kernel is based on a row-wise product method called SMASH [156]. Our method exploits high data reuse behavior [148] and minimizes the number of input matrix reads, while still maintaining low on-chip memory usage. SMASH incorporates on-chip memory to store intermediate results and leverages the atomic instructions to accumulate these partial products.
In this dissertation , we present a set of successive improvements, resulting in three versions of SMASH [156] (overview of SMASH architecture shown in Figure 7). In each version we identify the remaining bottlenecks, and then optimize our algorithm to mitigate them in the next version. Each SMASH implementation targets a specific performance bottleneck on the custom accelerator architecture.
@@
The following subsections describe our implementation of SMASH, along with three different optimizations. Similar to Gustavson’s two-phase matrix multiplication algorithm [73], our SMASH implementation is characterized by two phases:
@@ 1 Memory computation phase
Analogous to Gustavson’s two-phase matrix multiplication approach [73], the first phase of SMASH determines the memory required for the output matrix $C$, as well as the on-chip memory requirements of the intermediate products. For evaluation purposes, we do not include the time consumed in the memory computation phase while considering speedup over other architectures. Evaluation methodologies are further discussed in Section 5. This phase can be further decomposed into three tasks.
@@
    Compute Memory Requirements - After reading the input matrix arrays in CSR format, we compute the required size of memory required to store the output matrix by counting the total FMA operations per row. We compute the maximum number of non-zeros for every row of output matrix $C$ using Gustavson’s two-step algorithm [73]. In this dissertation , we refer to this term as Floating-Point Multiply-Add (FMA). The computation of FMA s per row has a computational complexity of $\mathcal{O}(n)$, where $n$ is the size of the input matrix.
@@ 2 Product Computation Phase
1 // READ PHASE
2 while Till you reach end of window do
3       // Atomically distribute work to each thread
4       $token\leftarrow$ Each thread will receive one unique token
5       if $token\_id\ \%\ 2=0$ then
@@ 9             $row\_begin\leftarrow A\_col\_ptr\_copy\_2[\frac{token\_id}{2}]$
12      $row\_end\leftarrow A\_col\_ptr[\frac{token\_id}{2}+1]$
13       for $i\leftarrow$ Iterate from $row\_begin$ to $row\_end$ do
@@ 12      $row\_end\leftarrow A\_col\_ptr[\frac{token\_id}{2}+1]$
15                   $col\_begin\leftarrow B\_row\_ptr[\frac{token\_id}{2}]$
16                   $col\_end\leftarrow B\_row\_ptr[\frac{token\_id}{2}+1]$
17                   if $token\_id\ \%\ 2=0$ then
@@ 16                   $col\_end\leftarrow B\_row\_ptr[\frac{token\_id}{2}+1]$
2       // Multiply element from $mat\_A$ with that from $mat\_B$ and store its tag and value
3       $tag\leftarrow$ $X$ coordinate from $mat\_A$ element and $Y$ coordinate from $mat\_B$ element
4       // Hash the Tag
5       $tag\leftarrow tag\ \%\ prime\_modulo$
6       if $SPAD\_tag[tag]=EMPTY$ then
7             $SPAD\_tag[tag]\leftarrow tag$ // Store Tag on scratchpad
8             $SPAD\_val[tag]\leftarrow value$ // Store Value on scratchpad
9
@@ 5 SMASH Evaluation
We evaluate the performance of our SMASH SpGEMM kernel implementation on synthetic, as well as real-world, datasets. For synthetic datasets, we chose the RMAT, as it produces datasets possessing a non-zero power-law distribution [33, 169], making it harder to find patterns in the non-zero values.
@@ 6 SMASH Summary
Chapter 5 Hardware Acceleration of GNNs
Graph Neural Networks (GNNs) are emerging as a formidable tool for processing non-euclidean data across various domains, including bioinformatics, financial networks, energy networks, telecommunication and social network analysis. Despite their effectiveness, their adoption has not been pervasive because of scalability challenges associated with large-scale graph datasets, particularly when leveraging message passing, posing significant computational bottlenecks. This class of large-scale workloads exhibits irregular sparsity patterns, resulting in unbalanced utilization of computational resources.
1 Key Bottlenecks in Accelerating GNNs
Deep Neural Networks have proven to be powerful models for solving problems that rely on data with an underlying Euclidean or grid-like structure [207], such as computer vision, natural language processing, and audio vision. In contrast, Graph Neural Networks (GNNs) have emerged as powerful frameworks for handling non-Euclidean data (e.g., social networks on the scale of billions [164]), achieving impressive performance across various domains such as social science, chemistry, and bioinformatics [197]. However, the computational complexity of GNNs, especially when working with ultra-sparse, large-scale graph datasets, poses challenges due to architectural limitations of traditional hardware (i.e., CPUs / GPUs) [146]. Moreover, GNNs predominantly adopt a recursive neighborhood aggregation methodology, in which each node aggregates the feature vectors of its neighboring nodes to derive its own updated feature vector. The scalability of message passing in GNNs, when applied to large graph structures, poses a significant bottleneck, especially as the size of the graphs surpasses the capacity of on-chip memory hierarchies in today’s CPUs and GPUs [159]. This leads to redundant and time-consuming memory transactions to fetch data from the main memory [14].
We identify the following three key bottlenecks for GNN workloads:
@@ 1 Key Bottlenecks in Accelerating GNNs
    Diverse Data Dependence Patterns: GNN workloads feature multiplication and accumulation operations, each demonstrating unique data dependency patterns.
@@
    Poor Hardware Resource Utilization: Compute units suffer from data starvation and load imbalance due to irregular sparsity patterns exhibited by large input graphs.